# A compact model of Escherichia coli core and biosynthetic metabolism

### v2

@@ 4.9 Enumeration of Elementary Flux Modes
Elementary Flux Modes (EFMs) enumeriation for the submodel iCH360red was performed, for each growth condition, using EFMtools [terzer_large-scale_2008]. Filtered modes (Supplementary Table 4) were defined as those supporting nonzero biomass flux and, for aerobic modes, nonzero oxygen uptake. In addition, in the aerobic case, filtered modes exclude those carrying flux in either of three reactions – Pyruvate-Formate Lyase (PFL), Fumarate Reductase (FRD2), and the menaquinone-dependent Dihydroorotate dehydrogenase (DHORD5) – known to be only physiologically active under anaerobic conditions [zhang_inactivation_2001, cecchini_succinate_2002, andrews_anaerobic_1977]
4.10 Growth/yield tradeoff analysis
To analyse trade-offs between growth rate and yield, we computed the yield of each EFM as the ratio of its biomass flux $v_{\rm BM}$ to its glucose uptake flux. In FBA models, fluxes are given in units of mmol/gDW/h, except for the biomass flux, whose conventional unit is $\rm{h}^{-1}$ because the biomass flux in these models is scaled to describe the biomass production rate per amount of biomass present. This means that our yields are given in units of gDW/mmol. To determine the cell growth rate allowed by a flux distribution, we first computed its absolute enzyme cost
@@ 4.10 Growth/yield tradeoff analysis
where $c_{\rm enz}$ is the enzyme cost of the flux distribution (an enzyme mass, measured in g/gDW), $v_{i}$ is the flux of reaction $i$, and $a_{i}$ is the enzyme cost per unit flux in reaction $i$, computed as per equation 1 using the set of adjusted turnover numbers. Clearly, both $v_{\rm BM}$ and $c_{\rm enz}$ depend on the particular choice of scaling of the mode (though their ratio, $v_{\rm BM}/c_{\rm enz}$, doesn’t). In order to obtain an estimate for the growth rate that is scale-invariant, we thus rescaled all modes to the same enzyme cost $f$ (in g/gDW), and looked at the resulting flux through the biomass reaction. More formally, the achievable growth rate $\mu$ (in $\rm h^{-1}$) for an (arbitrarily scaled) EFM with biomass flux $v_{\rm BM}$ and enzyme cost $c_{\rm enz}$ was computed as:
$\mu=f~{}\frac{v_{\rm BM}}{c_{\rm enz}}$ (5)
For simplicity, we approximate $f$ by a constant value of $f=0.285$ g/gDW, which we obtained by taking the minimum enzyme investment required by the enzyme-constrained model to support the experimentally measured growth rate reported in the proteomic dataset [schmidt_quantitative_2016] for the condition of interest in this analysis (aerobic growth on glucose).
To simulate low-oxygen conditions, the cost per unit flux of all oxygen-consuming reactions (CYTBO3_4pp, CYTBDpp, CYTBD2pp) was increased by a $1000$-fold, mimicking the physiological state in which these reaction operate at low substrate saturation.
@@ 4.11 Saturation FBA analysis
Saturation FBA calculations were performed by optimising biomass production in the enzyme-constrained model, setting the saturation coefficient ($\sigma$ in Eq. (1) to the value fitted as part of the turnover number adjusting procedure (Section 4.7) for all reactions except the glucose transporter (GLCptspp). The saturation of the glucose transporter, $\sigma_{\rm up}$, was computed as a function of external glucose concentration assuming irreversible Michaelis Mentens kinetics, so that:
@@
where $\rm[Glc]$ is the external glucose concentration (in mM) and a value of $0.116$ mM was used for the Michaelis constant $K_{\rm m}$ [wortel_metabolic_2018]
@@ 4.12 Component contribution estimates of thermodynamic constants
Estimates of the free energies of reactions, and their uncertainties, were obtained using the component contribution framework previously described in [noor_consistent_2013]. Several reactions in the model involve protein side groups, such as the acyl-carrying protein (ACP), or cofactors, such as glutaredoxin, for which a decomposition in terms of chemical groups is not available. As a result, thermodynamic constants for these reactions cannot be directly estimated through database-based implementations of the component contribution method, such as eQuilibrator [beber_equilibrator_2022]. However, since these non-decomposable protein groups are conserved in all reactions within our model, their net contribution to the reaction thermodynamics is, at least from a group contribution perspective, null. If we were only interested in computing standard free energies of reaction, $\Delta_{r}G^{\circ}$, we could simply treat protein groups as non-decomposable ”black-box” units and add them to the group incidence matrix of the eQuilibrator database (see Section S1.1 in [beber_equilibrator_2022]), enabling us to construct a group decomposition for compounds that contain them. However, for the computation of transformed standard free energies of reaction, $\Delta_{r}^{\prime}G^{\circ}$, an exact chemical definition of each metabolite is required.
To address this issue, protein groups were replaced by an appropriate chemical moiety that best approximates the metabolite’s chemical environment. Specifically, ACP- groups were replaced by a phosphopantetheine group, the natural prosthetic group of acyl-carrier proteins, with a methyl group at the attachment site to the protein scaffold. Similarly, glutaredoxin was replaced by its Cys-Pro-Tyr-Cys active site, with the two cysteines being either free (for the reduced form of the cofactor) or linked by a disulfide bridge (for the oxidised state). International Chemical Identifiers (InChI) were constructed for these ”replacement” metabolites and used to extend the default compound cache of the EQuilibrator database. The chemical identifiers for these compounds and the custom-extended EQuilibrator compound cache are available in the repository supporting this manuscript.
Corrections for reactions occurring across different compartments were computed as described in [beber_equilibrator_2022], using compartment-specific pH, pMg, ionic strength and potentials from [gollub_enkie_2023].
@@ 4.13 Max-min driving force computation
The max-min driving force (MDF) of each reference flux distribution was computed by extending the original formulation described in [noor_pathway_2014] to account for correlated uncertainty in the estimates of the thermodynamic constants. Let $\Delta G_{r}^{\prime\circ}\in\mathbb{R}^{N}$ be a random vector representing the (uncertain) standard Gibbs free energy of reaction for the reactions in the network. Importantly, this vector includes only balanced metabolic reactions and excludes pseudoreactions such as exchange reactions. To describe our uncertain knowledge, we assume that the vector $\Delta G_{r}^{\prime\circ}$ follows a multivariate normal distribution:
$\Delta G_{r}^{\prime\circ}\sim\mathcal{N}(\bar{\Delta G_{r}^{\prime\circ}},\Sigma)$ (7)
where $\bar{\Delta G_{r}^{\prime\circ}}$ and $\Sigma$ are the mean and covariance of the estimates obtained through the component contribution methods. This random vector can equivalently be expressed as:
$\Delta G_{r}^{\prime\circ}=\bar{\Delta G_{r}^{\prime\circ}}+\mathbf{Q}~{}\mathbf{z}$ (8)
where $\mathbf{z}$ is a standard normal random vector in $\mathbb{R}^{q}$, with $q=\text{rank}(\Sigma)$, and $\mathbf{Q}\in\mathbb{R}^{N\times q}$ is a square root of the covariance matrix, i.e. it satisfies:
$\Sigma=\rm\mathbf{Q}~{}\mathbf{Q}^{T}.$ (9)
@@
where $\mathbf{m}\in\mathbb{R}^{q}$ is a free parameter vector and $\chi_{q}(.)$ denotes the quantile function (inverse cumulative distribution function) of a chi-squared distribution with $q$ degrees of freedom. We can thus account for the uncertainty in thermodynamic estimates by treating the free energies of reaction as decision variables, rather than known parameters, and constraining their value to belong to ($\mathcal{D}_{\alpha}$). For a given reference flux distribution $\mathbf{v}$, this results in the following quadratically-constrained program (QCP):
$\underset{\mathbf{m},b}{\text{max}}$ $b$ (11)
s.t. $\Delta_{r}G^{\prime o}=\bar{{\Delta_{r}G^{\prime o}}}+\mathbf{Q}\mathbf{m}$
$\Delta_{r}G^{\prime}=\Delta_{r}G^{\prime o}+\rm{RT}~{}\mathbf{S}^{\top}\mathbf{c}$
@@
$\mathbf{l_{c}}\leq\mathbf{c}\leq\mathbf{u_{c}}$
@@ 4.14 Probabilistic thermodynamic analysis
Probabilistic metabolic optimisation (PMO) of the model was performed using the PTA python package [gollub_probabilistic_2021], providing the software with the curated thermodynamic estimates generated for this model. The default values available through the package for growth on M9 medium with glucose (which include measurements from [gerosa_pseudo-transition_2015]) were used as priors for the concentration of metabolites, and the growth rate was bounded below by the reported value in [gerosa_pseudo-transition_2015]. Further, for this analysis, which requires every flux in the solution to have a well-defined directionality, two transhydrogenase reactions (NADH17pp, and THD2pp) were allowed to operate reversibly [gollub_probabilistic_2021].
@@
$\eta=\frac{e^{-\frac{\Delta_{r}G^{\prime}}{\rm RT}}-1}{e^{-\frac{\Delta_{r}G^{\prime}}{\rm RT}}+1}$ (12)
where $\Delta_{r}G^{\prime}$ is the Gibbs free energy of reaction in the thermodynamic state, ${\rm R}$ is the universal gas constant and ${\rm T}$ is the temperature considered for the analysis ($310.15~{}K$). To compare the PTA-predicted flux force efficacies with experimental measurements of enzyme abundance, we selected reactions with carrying at least $2.5\%$ of the glucose uptake flux, and pooled these into two groups, corresponding to $\eta>0.5$ ($72$ reactions) and $\eta<0.5$ ($31$ reactions). $\eta=0.5$ corresponds to $\Delta_{r}G^{\prime}\approx-1.1~{}{\rm RT}=-2.8$ kJ/mol.
[…]
Figure 8: The flux force efficacy ($\eta$) as a function of the (scaled) negative Gibbs free energy of reaction, $-\frac{\Delta_{r}G^{\prime}}{\rm RT}$. The flux force efficacy corresponds to the ratio between net flux (forward minus backward flux) and total flux (forward plus backward flux) of a reaction, which approaches $1$ for reactions operating far from chemical equilibrium ($\Delta_{r}G^{\prime}\ll 0$).
@@ 5 Author contributions
A.B.-E. conceived the project and supervised its initial stage. M.C and H.H. assembled and curated the model. M.C. wrote software, performed simulations and analysed results. W.L., E.N. and H.H. supervised the project and contributed to the analysis and interpretation of results. M.C., H.H., W.L., and E.N. wrote the manuscript.
@@ 6 Competing interests
The authors declare that they comply with the PCI rule of having no financial conflicts of interest in relation to the content of the article.
@@ 7 Code availability
The model, together with all relevant data, is available on Github at https://github.com/marco-corrao/iCH360. The code and files required to reproduce all analysis in this manuscript are available on Github at https://github.com/marco-corrao/iCH360_paper and on Zenodo at https://doi.org/10.5281/zenodo.11092781.
@@ A.1 Changes to reactions from the model iML1515
After assembling the model as a subset of reactions from the genome-scale model iML1515, a few minor corrections were applied to some of the reactions based on evidence gathered from the literature. Note, however, that these changes did not result from an exhaustive process of review of the parent model. Such corrections include:
@@ A.2 Computation of an equivalent biomass reaction
In constraint-based models of metabolism, the biomass reaction summarises the production of all molecules that are not explicitly described by the model. These are typically macromolecules such as proteins or polynucleotides. Which metabolites are drained from the model network towards these other parts of metabolism, and in what proportions, depends on what compounds are described by the model. Hence, starting from an existing model, taking only a subset of the reactions, but leaving the biomass reaction unchanged would lead to inconsistencies.
@@
To construct a biomass reaction for iCH360 that corresponds equivalently to the biomass reaction in iML1515, the following method was used. First, we collected all the pathways required for the production of any component present in the iML1515 biomass reaction, but not in our model. These additional pathways (available in the repository supporting this manuscript) were manually curated based on available literature and database annotation to ensure they represent the most biologically-relevant bioproduction route for each biomass component. By adding these pathways to iCH360, we obtained an extended model (iCH360ext), which was able to grow directly through the original iML1515 biomass reaction. Let $N$ and $M$ denote the number of reactions in iCH360 and iCH360ext, respectively. Next, a reference flux distribution $\mathbf{v}^{*}_{ref}\in\mathbb{R}^{M}$ is computed on the extended model through FBA. We can express this flux vector as:
$\mathbf{v}^{*}_{ref}=\begin{pmatrix}\mathbf{v}^{*}\\ \mathbf{v}^{*}_{+}\\ v_{\rm BM}^{*}\end{pmatrix}$ (13)
@@
where $\mathbf{v}^{*}$ is the subset of its fluxes corresponding to reactions present in iCH360, $\mathbf{v}^{*}_{+}$ is the subset of fluxes corresponding to the reactions added to create the extended model iCH360ext, and $v_{\rm BM}^{*}$ is the flux through the biomass reaction. If the fluxes $\mathbf{v}^{*}$ were to be imposed on iCH360, a number of metabolites would necessarily remain unbalanced, that is:
${\rm\mathbf{S}}\mathbf{v}^{*}\neq\mathbf{0}$ (14)
@@
where ${\rm\mathbf{S}}$ denotes the stoichiometric matrix of iCH360. Hence, we compute the equivalent biomass reaction as the additional reaction (column of ${\rm\mathbf{S}}$) required to balance the system while achieving the same biomass flux of the reference distribution. That is:
$\Big{[}{\rm\mathbf{S}}|r_{eq}\Big{]}*\begin{pmatrix}\mathbf{v}^{*}\\ v_{\rm BM}^{*}\end{pmatrix}=\mathbf{0}$ (15)
which implies:
$r_{eq}=-\frac{1}{v_{\rm BM}^{*}}{\rm\mathbf{S}}\mathbf{v}^{*}$ (16)
@@
The stoichiometry of this equivalent biomass reaction is dependent on the choice of reference flux distribution $\mathbf{v}^{*}_{ref}$ and, generally, may be different for different growth conditions. This results from the fact that, depending on the condition, the extended model may produce the same biomass component through different pathways, which would then be converted by our procedure into different equivalent costs of precursors in the sub-model. Nevertheless, the additional biosynthesis pathways we used to form the extended model do not allow for alternative routes to biomass, hence we found the equivalent biomass reaction to be, in this case, unique across conditions.
A.3 Recursive computation of attributes in the annotation graph

### v3

@@ 4.10 Enumeration of Elementary Flux Modes
Elementary Flux Modes (EFMs) for the submodel iCH360red were enumerated, for each growth condition, using EFMtools [terzer_large-scale_2008]. Filtered modes (Supplementary Table 5) were defined as those supporting nonzero biomass flux and, for aerobic modes, nonzero oxygen uptake. In addition, in the aerobic case, filtered modes exclude those carrying flux in either of three reactions – Pyruvate-Formate Lyase (PFL), Fumarate Reductase (FRD2), and the menaquinone-dependent Dihydroorotate dehydrogenase (DHORD5) – known to be physiologically active only under anaerobic conditions [zhang_inactivation_2001, cecchini_succinate_2002, andrews_anaerobic_1977].
4.11 Growth/yield trade-off analysis
To analyse trade-offs between growth rate and yield, we computed the yield and a predicted growth rate of each EFM, using the unit conventions common in stoichiometric metabolic models. The specific growth rate $\mu$ (1/h) of a cell or a cell population can be defined as the rate of biomass production per amount of biomass present. In stoichiometric metabolic models, units are chosen in a specific way such that the growth rate is identically given by the rate $v_{\rm BM}$ of the biomass reaction. In FBA models, normal reaction fluxes are given in units of mmol/gDW/h, so the yield of an EFM, computed as the ratio of its biomass flux $v_{\rm BM}$ to its glucose uptake flux, has a unit of gDW/mmol. The biomass flux, whose conventional unit is $\rm{h}^{-1}$ was directly interpreted as the cell growth rate $\mu$. To determine the cell growth rate allowed by a flux distribution, we first computed its absolute enzyme cost
@@ 4.11 Growth/yield trade-off analysis
where $c_{\rm enz}$ is the enzyme cost of the flux distribution (an enzyme mass, measured in g/gDW), $v_{i}$ is the flux of reaction $i$, and $a_{i}$ is the enzyme cost per unit flux in reaction $i$, computed as per equation (1) using the set of adjusted turnover numbers. Since an elementary flux mode can be scaled arbitrarily, both $v_{\rm BM}$ and $c_{\rm enz}$ depend on the particular choice of scaling of the mode (though their ratio, $v_{\rm BM}/c_{\rm enz}$, does not). In order to obtain an estimate for the achievable growth rate of a mode $\mu$ that is a unique property of each EFM, independent of its scaling, we thus normalised all modes to the same total enzyme cost $f_{\rm enz}$ (in g/gDW), and looked at the resulting flux through the biomass reaction. More formally, the achievable growth rate $\mu$ (in $\rm h^{-1}$) for an (arbitrarily scaled) EFM with biomass flux $v_{\rm BM}$ and enzyme cost $c_{\rm enz}$ was computed as:
$\mu=f_{\rm enz}~{}\frac{v_{\rm BM}}{c_{\rm enz}}$ (5)
where $f_{\rm enz}$ denotes the total mass of enzyme available for the flux distribution, relative to the total dry mass of the cell. For simplicity, we approximate $f_{\rm enz}$ by a constant value of $f_{\rm enz}=0.285$ g/gDW, which we obtained by taking the minimum enzyme investment required by the enzyme-constrained model to support the experimentally measured growth rate reported in the proteomic data set [schmidt_quantitative_2016] for the condition of interest in this analysis (aerobic growth on glucose).
To simulate low-oxygen conditions, the cost per unit flux of all oxygen-consuming reactions (CYTBO3_4pp, CYTBDpp, CYTBD2pp) was increased by a $1000$-fold, mimicking the physiological state in which these reactions operate at low saturation with oxygen.
@@ 4.12 Saturation FBA analysis
Saturation FBA calculations were performed by optimising biomass production in the enzyme-constrained model, setting the saturation coefficient ($\sigma$ in Eq. (1) to the value fitted as part of the turnover number adjusting procedure (Section 4.8) for all reactions except the glucose transporter (GLCptspp). The saturation of the glucose transporter, $\sigma_{\rm up}$, was computed as a function of external glucose concentration and assuming irreversible Michaelis Mentens kinetics, so that:
@@
where $\rm[Glc]$ is the external glucose concentration (in mM) and a value of $0.116$ mM was used for the Michaelis constant $K_{\rm m}$ [wortel_metabolic_2018].
@@ 4.13 Component contribution estimates of thermodynamic constants
Estimates of the free energies of reactions, and their uncertainties, were obtained using the component contribution framework previously described in [noor_consistent_2013]. Several reactions in the model involve protein side groups, such as the acyl-carrying protein (ACP), or cofactors, such as glutaredoxin, for which a decomposition in terms of chemical groups is not available. As a result, thermodynamic constants for these reactions cannot be directly estimated through database-based implementations of the component contribution method, such as eQuilibrator [beber_equilibrator_2022]. However, since these non-decomposable protein groups are conserved in all reactions within our model, their net contribution to the reaction thermodynamics is, at least from a group contribution perspective, null. If we were only interested in computing standard free energies of reaction, $\Delta_{r}G^{\circ}$, we could simply treat protein groups as non-decomposable “black-box” units and add them to the group incidence matrix of the eQuilibrator database (see Section S1.1 in [beber_equilibrator_2022]), enabling us to construct a group decomposition for compounds that contain them. However, for the computation of transformed standard free energies of reaction, $\Delta_{r}G^{\prime\circ}$, an exact chemical definition of each metabolite is required.
To address this issue, protein groups were replaced by an appropriate chemical moiety that best approximates the metabolite’s chemical environment. Specifically, ACP- groups were replaced by a phosphopantetheine group, the natural prosthetic group of acyl-carrier proteins, with a methyl group at the attachment site to the protein scaffold. Similarly, glutaredoxin was replaced by its Cys-Pro-Tyr-Cys active site, with the two cysteines being either free (for the reduced form of the cofactor) or linked by a disulfide bridge (for the oxidised state). International Chemical Identifiers (InChI) were constructed for these “replacement” metabolites (Supplementary File S3) and used to extend the default compound cache of the eQuilibrator database. This custom-extended eQuilibrator compound cache is available in the repository supporting this manuscript.
Corrections for reactions that occur in different compartments were calculated as described in [beber_equilibrator_2022], using compartment-specific pH, pMg, ionic strength, and potentials from [gollub_enkie_2023].
@@ 4.14 Max-min driving force computation
The max-min driving force (MDF) of each reference flux distribution was calculated by extending the original formulation described in [noor_pathway_2014] to account for correlated uncertainty in the estimates of the thermodynamic constants. Let $\Delta G_{r}^{\prime\circ}\in\mathbb{R}^{N}$ be a random vector representing the (uncertain) standard Gibbs free energy of reaction for the reactions in the network. Importantly, this vector includes only balanced metabolic reactions and excludes pseudoreactions such as exchange reactions. To describe our uncertain knowledge, we assume that the vector $\Delta G_{r}^{\prime\circ}$ follows a multivariate normal distribution:
$\Delta G_{r}^{\prime\circ}\sim\mathcal{N}(\Delta\bar{G}_{r}^{\prime\circ},\bm{\Sigma})$ (7)
where $\Delta\bar{G}_{r}^{\prime\circ}$ and ${\bm{\Sigma}}$ are the mean vector and covariance matrix of the estimates obtained through the component contribution methods. This random vector can equivalently be expressed as:
$\Delta G_{r}^{\prime\circ}=\Delta\bar{G}_{r}^{\prime\circ}+\mathbf{Q}~{}\mathbf{z}$ (8)
where $\mathbf{z}$ is a standard normal random vector in $\mathbb{R}^{q}$, with $q=\text{rank}(\bm{\Sigma})$, and $\mathbf{Q}\in\mathbb{R}^{N\times q}$ is a square root of the covariance matrix, i.e. it satisfies:
$\bm{\Sigma}=\rm\mathbf{Q}~{}\mathbf{Q}^{T}.$ (9)
@@
where $\mathbf{m}\in\mathbb{R}^{q}$ is a vector of free parameters and $\chi_{q}(.)$ denotes the quantile function (inverse cumulative distribution function) of a chi-squared distribution with $q$ degrees of freedom. We can thus account for the uncertainty in thermodynamic estimates by treating the free energies of reaction as decision variables, rather than known parameters, and constraining their value to belong to $\mathcal{D}_{\alpha}$. For a given reference flux distribution $\mathbf{v}$, this results in the following quadratically-constrained program (QCP):
$\underset{\mathbf{m},\mathbf{c},b}{\text{max}}$ $b$ (11)
s.t. $\Delta_{r}G^{\prime\circ}=\Delta\bar{G}_{r}^{\prime\circ}+\mathbf{Q}~{}\mathbf{m}$
$\Delta_{r}G^{\prime}=\Delta_{r}G^{\prime\circ}+\rm{RT}~{}\mathbf{S}^{\top}\mathbf{c}$
@@
$\mathbf{c_{\rm min}}\leq\mathbf{c}\leq\mathbf{c_{\rm max}}$
@@ 4.15 Probabilistic thermodynamic analysis
Probabilistic metabolic optimisation (PMO) of the model was performed using the PTA Python package [gollub_probabilistic_2021], providing the software with the curated thermodynamic estimates generated for this model. The default values available through the package for growth in M9 medium with glucose (which include measurements from [gerosa_pseudo-transition_2015]) were used as priors for the concentration of metabolites, and the growth rate was bounded from below by the reported value in [gerosa_pseudo-transition_2015]. Furthermore, for this analysis, which requires that each flux in the solution have a well-defined directionality, two transhydrogenase reactions (NADH17pp and THD2pp) were allowed to operate reversibly [gollub_probabilistic_2021].
@@
$\eta=\frac{\mathrm{e}^{-\frac{\Delta_{r}G^{\prime}}{\rm RT}}-1}{\mathrm{e}^{-\frac{\Delta_{r}G^{\prime}}{\rm RT}}+1}=\tanh\Big{(}-\frac{1}{2}\frac{\Delta_{r}G^{\prime}}{\rm RT}\Big{)}$ (12)
where $\Delta_{r}G^{\prime}$ is the Gibbs free energy of reaction in the thermodynamic state, ${\rm R}$ is the universal gas constant and ${\rm T}$ is the temperature considered for the analysis ($310.15~{}K$). To compare the PTA-predicted flux force efficacies with experimental measurements of enzyme abundance, we selected reactions that carried at least $2.5\%$ of the glucose uptake flux and pooled them into two groups, corresponding to $\eta>0.5$ ($72$ reactions) and $\eta<0.5$ ($31$ reactions). $\eta=0.5$ corresponds to $\Delta_{r}G^{\prime}\approx-1.1~{}{\rm RT}=-2.8$ kJ/mol.
00222244446666000.50.50.50.51111Driving Force ($-\frac{\Delta_{r}G^{\prime}}{\rm RT}$)Thermodynamic efficiency ($\eta$)
Figure 8: The flux force efficacy ($\eta$) as a function of the (scaled) negative Gibbs free energy of reaction, $-\frac{\Delta_{r}G^{\prime}}{\rm RT}$. The efficacy of the flux force corresponds to the ratio between the net flux (forward minus backward flux) and the total flux (forward plus backward flux) of a reaction, which approaches $1$ for reactions operating far from chemical equilibrium ($\Delta_{r}G^{\prime}\ll 0$).
@@ 5 Author contributions
A.B.-E. conceived the project and oversaw its initial stage. M.C and H.H. assembled and curated the model. M.C. wrote software, performed simulations and analysed results. W.L., E.N. and H.H. oversaw the project and contributed to the analysis and interpretation of results. M.C., H.H., W.L., and E.N. wrote the manuscript.
@@ 6 Competing interests
The authors declare that they comply with the PCI rule of having no financial conflicts of interest with respect to the content of the article.
@@ 7 Code availability
The model, together with all relevant data, is available on Github at https://github.com/marco-corrao/iCH360. The code and files required to reproduce all analyses in this manuscript are available on Github at https://github.com/marco-corrao/iCH360_paper and on Zenodo at https://doi.org/10.5281/zenodo.11092781. Additional analyses and comparisons with other models are also available in these repositories.
@@ A.1 Changes to reactions from the model iML1515
After assembling the model as a subset of reactions from the genome-scale model iML1515, a few minor corrections were applied to some of the reactions based on evidence gathered from the literature. Note, however, that these changes did not result from an exhaustive process of review of the parent model. The corrections include:
@@ A.2 Computation of an equivalent biomass reaction
In constraint-based models of metabolism, the biomass reaction summarises the production of all molecules that are not explicitly described by the model. These are typically macromolecules such as proteins or polynucleotides. Which metabolites are drained from the model network towards these other parts of metabolism, and in what proportions, depends on what compounds are described by the model. Hence, starting from an existing model, taking only a subset of the reactions but leaving the biomass reaction unchanged would lead to inconsistencies.
@@
To construct a biomass reaction for iCH360 that corresponds equivalently to the biomass reaction in iML1515, the following method was used. First, we collected all the pathways required for the production of the components present in the iML1515 biomass reaction (with the exception of a small number of compounds present with very small stoichiometry, which were excluded from this analysis for simplicity), but not in our model. These additional pathways (available in the repository supporting this manuscript) were manually curated based on available literature and database annotation to ensure they represent the most biologically-relevant bioproduction route for each biomass component. By adding these pathways to iCH360, we obtained an extended model (iCH360ext), which was able to predict growth directly through the original biomass reaction.
With this extended model at hand, we can continue as follows. Let $N$ and $M$ denote the number of reactions and metabolites, respectively, in iCH360. Further, let $N_{\rm ext}>N$ denote the number of reactions in the extended model iCH360ext. A reference flux distribution $\mathbf{v}^{*}_{\rm ref}\in\mathbb{R}^{N_{\rm ext}}$ is computed on the extended model through FBA. We can express this flux vector as:
$\mathbf{v}^{*}_{\rm ref}=\begin{pmatrix}\mathbf{v}^{*}\\ \mathbf{v}^{*}_{+}\\ v_{\rm BM}^{*}\end{pmatrix}$ (13)
@@
where $\mathbf{v}^{*}$ is the subset of its fluxes corresponding to reactions present in iCH360, $\mathbf{v}^{*}_{+}$ is the subset of fluxes corresponding to the reactions added to create the extended model iCH360ext, and $v_{\rm BM}^{*}$ is the flux through the iML1515 biomass reaction. If the fluxes $\mathbf{v}^{*}$ were to be imposed on iCH360 (which, at this stage, does not yet contain a biomass reaction), a number of metabolites would necessarily remain unbalanced, that is:
${\rm\mathbf{S}}~{}\mathbf{v}^{*}\neq\mathbf{0}$ (14)
@@
where ${\rm\mathbf{S}}\in\mathbb{R}^{M\times N}$ denotes the stoichiometric matrix of iCH360. We now seek to define an ”equivalent” biomass reaction that, when added to the iCH360 network, will drain or produce metabolites so as to balance equation (14). Let $\mathbf{r_{\rm eq}}\in\mathbb{R}^{M}$ denote the stoichiometry of such biomass reaction. We compute $\mathbf{r_{\rm eq}}$ as the additional column of ${\rm\mathbf{S}}$ required to balance the system while achieving the same biomass flux of the reference distribution. That is, $\mathbf{r_{\rm eq}}$ satisfies:
$\Big{(}{\rm\mathbf{S}}\;\mathbf{r_{\rm eq}}\Big{)}\begin{pmatrix}\mathbf{v}^{*}\\ v_{\rm BM}^{*}\end{pmatrix}=\mathbf{0}$ (15)
which can be solved as:
$\mathbf{r_{\rm eq}}=-\frac{1}{v_{\rm BM}^{*}}{\rm\mathbf{S}}\mathbf{v}^{*}$ (16)
@@
The stoichiometry of this equivalent biomass reaction depends on the choice of reference flux distribution $\mathbf{v}^{*}_{\rm ref}$ and, generally, may be different for different growth conditions. This results from the fact that, depending on the condition, the extended model may produce the same biomass component through different pathways, which would then be converted by our procedure into different equivalent costs of precursors in the sub-model. Nevertheless, the additional biosynthesis pathways we used to form the extended model do not allow for alternative routes to biomass, hence we found the equivalent biomass reaction to be, in this case, unique across conditions.
A.3 Computation of attributes using the knowledge graph